OpenStudy (anonymous):

The sum of two numbers is 50. One number is 10 less than three times the other number. What are the two numbers?

OpenStudy (anonymous):

A. 25 and 25B. 45 and 5C. 15 and 35 D. 40 and 10E. 20 and 30

OpenStudy (anonymous):

here is an example Let the larger number be x and the smaller number be y the sum of two numbers is twenty-one x + y = 21 y = 21 – x . . . . . . (1) twice the larger is 4 times the smaller number 2x = 4y x = 2y . . . . . . . . .(2) Substitute (1) into (2), x = 2(21 – x) x = 42 – 2x x + 2x = 42 3x = 42 x = 14 Substitute x = 14 into (1) y = 21 – 14 y = 7 Answer: x = 14, y = 7

OpenStudy (anonymous):

it makes sense but unfortunately 14 and 7 aren't part of the choices ):

OpenStudy (amistre64):

if it makes sense .... then you should adapt the example to your own specific problem

OpenStudy (anonymous):

im trying

OpenStudy (amistre64):

my idea would be to define the 2 numbers as: n is one of the numbers 3n-10 is the other number simply by the definition of the problem adding them has to be 50 so ... n + 3n - 10 = 50 once you solve for n you can then define the numbers

OpenStudy (anonymous):

that's a good way to do it to... were not just going to give you the answer .. then there is no point for this site and u might be better of looking up the answer on google

OpenStudy (anonymous):

15 & 35, thanks i got it